Choose the aqueous solution that has the highest boiling point.These are all solutions of nonvolatile solutes and you shouldassume ideal van't Hoff factors where applicable.

a. 0.100 m NaNO3

b. 0.100m Li2SO4

c. 0.200m C3H8O3

d. 0.060m Na3PO4
Answer #1
a)   molality of ions   =   0.2m         (Na+   and NO3-     so 2*0.1 = 0.2m)
b)   molality = 0.3     (2Li+   and 1 SO42-
c)   molaity = 0.2molal       (glycerol does notdissociate)
d)    molality = 0.06*4 =0.24     (3 Na+   and 1PO43-)


Highest molality of ions is Li2SO4. So b will have the highestboiling point

recall that boiling point elevation = Kb *van't HOff* molality ofsalt = Kb*molaity of ions

Kb = boiling point elevation for water

b) has highest molality of ions